### v1

@@ 2 Related Work
The prospect of models autonomously discovering effective reasoning strategies has motivated significant research in reinforcement learning approaches. Early work in teaching language models to reason with verifiable outcomes explored various RL methodologies, from off-policy and batch methods (Zelikman et al., 2022; Havrilla et al., 2024; Hoffman et al., 2023) to on-policy approaches (Zelikman et al., 2024; Kazemnejad et al., 2024; Cui et al., 2025). These methods differ in their approaches to credit assignment in reasoning trajectories. A notable breakthrough came with Deepseek-R1 (Guo et al., 2025; Jaech et al., 2024), which demonstrated that even a simplified version of PPO (Schulman et al., 2017) called GRPO (Shao et al., 2024) could lead to significant improvements and the emergence of in-context search behaviors. Recent analyses by Li et al. (2025), Liu et al. (2025), and Yeo et al. (2025) have begun to unpack these results, revealing that supervised fine-tuning with long, structured chains of thought enhances both the efficiency and performance of RL compared to shorter reasoning chains. However, a crucial question remains unanswered: why do some models successfully learn through RL while others fail to improve? Our work addresses this gap by investigating the essential properties of initial models that enable successful reinforcement learning of reasoning behaviors.
@@ 3.1 Initial Investigation: A tale of two models
[…]
@@ 3.2 A Framework for Analyzing Cognitive Behaviors
Each behavior can be identified by its pattern in the reasoning tokens. Backtracking is seen as token sequences that explicitly contradict and replace previous steps, verification produces tokens that compare outcomes against solution criteria, backward chaining generates tokens that construct solution paths from the goal to the initial state, and subgoal setting explicitly proposes an intermediate step to target on the path to the final goal. We develop a classification pipeline using GPT-4o-mini [^3] We chose GPT-4o-mini to balance inference cost and model capability to reliably identify these patterns in the outputs of the model.
@@ 3.3 The Role of Initial Behaviors in Self-Improvement
To better understand this disparity, we analyzed the baseline reasoning patterns across three models: Qwen-2.5-3B, Llama-3.2-3B, and Llama-3.1-70B. Our analysis revealed that Qwen-2.5-3B naturally exhibits substantially higher rates of all four behaviors compared to both Llama variants (Fig. 4). While the larger Llama-3.1-70B showed generally increased activation of these behaviors compared to Llama-3.2-3B, this improvement was notably uneven — backtracking, in particular, remained limited even in the larger model.
These observations suggest two insights: 1) certain cognitive behaviors in the initial policy may be necessary for models to effectively utilize increased test-time compute through extended reasoning sequences 2) increased model scale can improve the contextual activation of these behaviors. This pattern is particularly significant because reinforcement learning can only amplify behaviors that appear in successful trajectories — making these initial behavioral capabilities a precondition for effective learning.
@@ 3.4 Intervening on initial behaviors
We begin by curating seven distinct priming datasets using Countdown problems. Five of these datasets emphasize different behavioral combinations: all strategies combined, backtracking only, backtracking with verification, backtracking with subgoal setting, and backtracking with backward chaining. We generate these datasets using Claude-3.5-Sonnet [^4] We chose Claude-3.5-Sonnet as it reliably followed instructions to show the desired behaviors we wanted in the reasoning trajectories., leveraging its ability to produce reasoning trajectories with precisely specified behavioral characteristics. While Claude does not always produce the correct answer (see Fig. 9), it consistently demonstrates the requested reasoning patterns, providing clean behavioral primitives for our analysis.
To verify that improvement stems from specific cognitive behaviors rather than simply increased computation time, we introduce two control conditions: an empty chain-of-thought and a length-matched (to the datapoints from the all-strategies dataset) chain filled with placeholder tokens. These control datasets help us verify whether any observed improvements stem from specific cognitive behaviors rather than simply increased computation time. We also create a variant of our all-strategies dataset containing only incorrect solutions while maintaining the desired reasoning patterns. This variant allows us to disentangle the importance of cognitive behaviors from the accuracy of solutions.
@@
Surprisingly, models primed with incorrect solutions, but with the right behaviors achieve identical performance to those trained on datasets with correct solutions (Fig. 6). This suggests that the presence of cognitive behaviors, rather than the access to correct solutions, is the crucial factor enabling successful self-improvement through reinforcement learning. This extends prior work demonstrating learning from corrupted reasoning trajectories (Li et al., 2025) in a significant way — we show that reasoning patterns from a weaker models can effectively bootstrap the learning process to build more capable ones, suggesting that the presence of cognitive behaviors matter more than just the correctness of the outcomes.
@@ 3.5 Selectively amplifying behaviors in pretraining data
Behavioral Frequencies in Pretraining Data.
We first analyze the natural frequency of cognitive behaviors in pre-training data, focusing on OpenWebMath (Paster et al., 2023), and FineMath (Allal et al., 2025), which are constructed specifically for mathematical reasoning. Using Qwen-2.5-32B as a classifier [^5] We use Qwen-2.5-32B to balance throughput, capability and cost, we analyze 200,000 randomly sampled documents for the presence of our target behaviors. Even in this math-focused corpus, cognitive behaviors such as backtracking and verifications appear infrequently, suggesting that standard pretraining provides limited exposure to these crucial patterns (see Fig. 7).
@@ Acknowledgments
We would like to thank Charlie Snell, Eric Zelikman, Archit Sharma, Rafael Rafailov, Alex Havrilla, Ulyana Piterburg, Chase Blagden and Dimitris Papailiopoulos for their discussions and support. KG was supported by an HAI-SAP Grant and an NSF Expeditions grant; AS was supported by the NSF GRFP Fellowship. Compute for this work was provided by Synthlabs, Stanford Marlowe and the Toyota Research Institute (TRI).
@@ Appendix B Priming
To control specific reasoning behaviors across each dataset, we implemented customized system prompts that explicitly guide Claude toward the desired reasoning patterns. These system prompts  [^6] All prompts are available in our Github repository: https://github.com/kanishkg/cognitive-behaviors. were carefully crafted to elicit distinct problem-solving approaches while maintaining consistency across datasets. To create datasets rich in only the targeted behaviors and determine causal efficacy, we instructed Claude to use exclusively the specified cognitive behavior while prohibiting all others. For example, in the "Backtracking Only" condition, we emphasized through the system prompt that Claude could only use backtracking and was explicitly not permitted to verify answers, set subgoals, or work backwards from the target. For each of the five datasets, we selected 1,200 unique Countdown games as our problem set. We prompted Claude to solve these games while adhering to the specified reasoning strategy for each dataset. We then evaluated the accuracy of Claude’s solution trajectories, with results presented in Fig. 9. To prepare for model training, we divided the 1,200 examples in each dataset into 1,000 training samples and 200 evaluation samples. We fine-tuned our different base models (Llama3.2-3B and Qwen2.5-3B) for 5 epochs using the AdamW optimizer (Loshchilov & Hutter, 2019). The optimizer was setup to have a peak learning rate of 1e-5 and linear warm-up for the first 5% steps, followed by cosine annealing.
@@ Appendix C Reinforcement Learning
Appendix D Metrics.

### v2

@@ 2 Related Work
The prospect of models autonomously discovering effective reasoning strategies has motivated significant research in reinforcement learning approaches. Early work in teaching language models to reason with verifiable outcomes explored various RL methodologies, from off-policy and batch methods (Zelikman et al., 2022; Havrilla et al., 2024; Hoffman et al., 2023) to on-policy approaches (Zelikman et al., 2024; Kazemnejad et al., 2024; Cui et al., 2025). These methods differ in their approaches to credit assignment in reasoning trajectories. A notable breakthrough came with Deepseek-R1 (Guo et al., 2025; Jaech et al., 2024), which demonstrated that even a simplified version of PPO (Schulman et al., 2017) called GRPO (Shao et al., 2024) could lead to significant improvements and the emergence of in-context search behaviors. Recent analyses (Li et al., 2025), Wu et al. (2024), Liu et al. (2025), Ye et al. (2025) and Yeo et al. (2025) have begun to unpack these results, revealing that supervised fine-tuning with long, structured chains of thought enhances both the efficiency and performance of RL compared to shorter reasoning chains. However, a crucial question remains unanswered: why do some models successfully learn through RL while others fail to improve? (Yeo et al., 2025)briefly discuss this issue by analyzing the frequency of phrases related to reflective reasoning in the pretraining dataset but their analysis remains limited. Our work addresses this gap by investigating the essential properties of initial models that enable successful reinforcement learning of reasoning behaviors.
@@ 3.1 Initial Investigation: A tale of two models
We begin by investigating a surprising observation: language models of comparable size but from different families show markedly different capacities for improvement through reinforcement learning. The Countdown game serves as our primary testbed — a mathematical puzzle where players must combine a set of input numbers using the four basic arithmetic operations ($+,-,\times,\div$) to reach a target number. For example, given the numbers 25, 30, 3, 4 and a target of 32, the solution involves combining these numbers through a series of operations to reach exactly 32: $(30-25+3)\times 4$.
@@ 3.2 A Framework for Analyzing Cognitive Behaviors
Each behavior can be identified by its pattern in the reasoning tokens. Backtracking is seen as token sequences that explicitly contradict and replace previous steps, verification produces tokens that compare outcomes against solution criteria, backward chaining generates tokens that construct solution paths from the goal to the initial state, and subgoal setting explicitly proposes an intermediate step to target on the path to the final goal. We develop a classification pipeline using GPT-4o-mini [^3] We chose GPT-4o-mini to balance inference cost and model capability to reliably identify these patterns in the outputs of the model (see App. F for analysis of agreement with humans and larger models).
@@ 3.3 The Role of Initial Behaviors in Self-Improvement
To better understand this disparity, we analyzed the baseline reasoning patterns across three models: Qwen-2.5-3B, Llama-3.2-3B, and Llama-3.1-70B. Our analysis revealed that Qwen-2.5-3B naturally exhibits substantially higher rates of all four behaviors compared to both Llama variants (Fig. 4). While the larger Llama-3.1-70B showed generally increased activation of these behaviors compared to Llama-3.2-3B, this improvement was notably uneven — backtracking, in particular, remained limited even in the larger model. These observations suggest two insights: 1) certain cognitive behaviors in the initial policy may be necessary for models to effectively utilize increased test-time compute through extended reasoning sequences 2) increased model scale can improve the contextual activation of these behaviors. This pattern is particularly significant because reinforcement learning can only amplify behaviors that appear in successful trajectories — making these initial behavioral capabilities a precondition for effective learning.
@@ 3.4 Intervening on initial behaviors
We begin by curating seven distinct priming datasets using Countdown problems. Five of these datasets emphasize different behavioral combinations: all strategies combined, backtracking only, backtracking with verification, backtracking with subgoal setting, and backtracking with backward chaining. We generate these datasets using Claude-3.5-Sonnet [^4] We chose Claude-3.5-Sonnet as it reliably followed instructions to show the desired behaviors we wanted in the reasoning trajectories., leveraging its ability to produce reasoning trajectories with precisely specified behavioral characteristics. While Claude does not always produce the correct answer (see Fig. 9), it consistently demonstrates the requested reasoning patterns, providing clean behavioral primitives for our analysis. To verify that improvement stems from specific cognitive behaviors rather than simply increased computation time, we introduce two control conditions: 1) an empty chain-of-thought, where the thoughts are empty, “<think></think>” and 2) a chain-of-thought where the thoughts have the same length as those of the all-strategies control, but filled with placeholder tokens, “<think>. . . . . </think>”. We also create a variant of our all-strategies dataset containing only incorrect solutions while maintaining the desired reasoning patterns. This variant allows us to disentangle the importance of cognitive behaviors from the accuracy of solutions.
@@
Surprisingly, models primed with incorrect solutions, but with the right behaviors achieve identical performance to those trained on datasets with correct solutions (Fig. 6). This suggests that the presence of cognitive behaviors in the priming data, rather than the access to correct solutions, is the crucial factor enabling successful self-improvement through reinforcement learning. This extends prior work demonstrating learning from corrupted reasoning trajectories (Li et al., 2025) in a significant way — we show that reasoning patterns from weaker models can effectively bootstrap the learning process to build more capable ones, suggesting that the presence of cognitive behaviors matter more than just the correctness of the outcomes.
Behavioral Frequencies in Pretraining Data.
We first analyze the natural frequency of cognitive behaviors in pre-training data, focusing on OpenWebMath (Paster et al., 2023), and FineMath (Allal et al., 2025), which are constructed specifically for mathematical reasoning. Using Qwen-2.5-32B as a classifier [^5] We use Qwen-2.5-32B to balance throughput, capability and cost., we analyze 200,000 randomly sampled documents for the presence of our target behaviors. Even in this math-focused corpus, cognitive behaviors such as backtracking and verifications appear infrequently, suggesting that standard pretraining provides limited exposure to these crucial patterns (see Fig. 7).
@@ Acknowledgments
We would like to thank Charlie Snell, Eric Zelikman, Archit Sharma, Rafael Rafailov, Alex Havrilla, Ulyana Piterburg, Chase Blagden, Dimitris Papailiopoulos, and Jan-Philipp Fränken for their discussions and support. KG was supported by an HAI-SAP Grant and an NSF Expeditions grant; AS was supported by the NSF GRFP Fellowship. Compute for this work was provided by Synthlabs, Stanford Marlowe and the Toyota Research Institute (TRI).
@@ Appendix B Priming
Figure 10: Behaviors present in priming datasets. Analyzing the average counts of behaviors when Claude is instructed to only show specific behaviors in its thoughts.
To control specific reasoning behaviors across each dataset, we implemented customized system prompts that explicitly guide Claude toward the desired reasoning patterns. These system prompts  [^6] All prompts are available in our Github repository: https://github.com/kanishkg/cognitive-behaviors. were carefully crafted to elicit distinct problem-solving approaches while maintaining consistency across datasets (see Fig. 10 for an analysis of behaviors present). To create datasets rich in only the targeted behaviors and determine causal efficacy, we instructed Claude to use exclusively the specified cognitive behavior while prohibiting all others. For example, in the "Backtracking Only" condition, we emphasized through the system prompt that Claude could only use backtracking and was explicitly not permitted to verify answers, set subgoals, or work backwards from the target. For each of the five datasets, we selected 1,200 unique Countdown games as our problem set. We prompted Claude to solve these games while adhering to the specified reasoning strategy for each dataset. We then evaluated the accuracy of Claude’s solution trajectories, with results presented in Fig. 9.
We provide the SFT hyperparameters in Tab. 1.
Data & Model
Training/Validation Dataset Size 1000 / 200
Context Window 2048
Training/Validation Batch Size 64 / 64
Optimization
Optimizer AdamW (Loshchilov & Hutter, 2019)
Peak Learning Rate 1e-5
Warmup Linear (5% of total steps)
Annealing Cosine
Total Epochs 5
Table 1: SFT Hyperparameters
@@ Appendix C Reinforcement Learning
Table 2: PPO Training Hyperparameters
Appendix D Metrics
@@
Appendix F Interrater Reliability for Classifying Behaviors
To assess the reliability of classification and counting of behaviors from GPT-4o-mini, we conducted an inter-rater reliability analysis using the Intraclass Correlation Coefficient (ICC3). We compared agreement of GPT-4o-mini, Claude, and two human annotators — of four key problem-solving behaviors: verification, backtracking, subgoal setting, and backward chaining. We randomly sample 100 reasoning trajectories from different conditions and trained models to measure agreement.
GPT-4o-mini showed (see Fig. 11) strong agreement with other raters across most behaviors, though with varying levels of consistency. For verification, GPT-4o-mini achieved substantial agreement with both Claude (ICC3 = 0.70) and human raters (ICC3 = 0.66 and 0.65). For backtracking, GPT-4o-mini maintained robust reliability with Claude (ICC3 = 0.74) and slightly lower but consistent agreement with human raters (ICC3 = 0.67). Subgoal setting showed the strongest consistency, with GPT-4o-mini achieving high agreement with Claude (ICC3 = 0.88) and human raters (ICC3 = 0.70 and 0.77). Backward chaining demonstrated lower agreement levels, with GPT-4o-mini showing stronger correlation with Claude (ICC3 = 0.78) than with human raters (ICC3 = 0.55 and 0.40). These results suggest that GPT-4o-mini can reliably identify and classify most cognitive behaviors.
Figure 11: Inter-rater reliability for counting and classifying behaviors. Intraclass Correlation Coefficient (ICC3) values for (a) Verification (b) Backtracking (c) Subgoal Setting (d) Backward chaining for labels from GPT-4o-mini, Claude, and two human raters.
Appendix G Test-time Scaling.
Figure 12: Test-time Scaling by Varying Token Budgets during Inference. The graph shows how different training checkpoints of the Llama model, fine-tuned with RL using the all-strategies primed model when tested with different maximum token lengths (128, 256, 512, and 1024 tokens). This analysis demonstrates how model performance scales with available computational resources at inference time.
We test how model performance scales with different token budget constraints at inference time. By varying the maximum allowed token length from $2^{7}$ (128) to $2^{10}$ (1024) tokens, we analyzed the relationship between computational resources and model performance. We test training checkpoints ranging from 0 to 200 steps of PPO training for the all-strategies primed Llama model. Performance consistently improves with larger token budgets across all checkpoints, but with diminishing returns after 512 tokens (see Fig. 12). Later checkpoints (150-200 steps) show stronger performance even with restricted token budgets, achieving scores of $\approx$0.4 with just 128 tokens compared to near-zero performance for early checkpoints under the same constraints. This suggests that training helps the model learn to use limited token budgets more efficiently — more train-time interactions lead to more efficient test-time performance (Jones, 2021). The scaling behavior also reveals an interesting convergence pattern - while early checkpoints (0-20 steps) show nearly log-linear scaling with token budget increases, later checkpoints exhibit more logarithmic scaling, suggesting they have learned to make better use of additional tokens. By 1024 tokens, the best checkpoints achieve scores around 0.7, with relatively small gaps between checkpoints after 100 training steps.
Appendix H Transfer of Behaviors to Other Domains.
To evaluate the generalization of learned cognitive behaviors, we investigated whether models trained to exhibit specific problem-solving strategies in mathematical reasoning could transfer these behaviors to other data distributions. First, we compared two variants: a model trained with a curated pretraining dataset designed to amplify cognitive behaviors in mathematical reasoning, and a control model trained on behavior-minimized data on 200 questions from GPQA (Google Proof Question Answering; Rein et al. (2024)). The results (see Fig. 13) demonstrate significant transfer of learned behaviors, with the curated pretraining model exhibiting substantially higher frequencies of all four cognitive strategies (verification, backtracking, subgoal setting, and backward chaining) when solving general knowledge questions. Most notably, subgoal setting showed the strongest transfer effect, with an average of 6.5 instances per question in the curated model compared to 0.7 in the control. This suggests that cognitive behaviors amplified in mathematical reasoning can generalize effectively to broader question-answering contexts. It should be noted that there is no significant performance difference between the two models (both score about $\approx 12$%). This performance parity despite behavioral differences suggests is because the model is limited in its forward inference capabilities rather than its problem-solving strategy — in other words, the model can learn to approach problems more systematically, but still struggles with the basic reasoning and knowledge needed to arrive at correct answers.
Figure 13: Transfer of Behaviors to Question Answering. Average frequency of four cognitive behaviors (verifications, backtracking, subgoal setting, and backward chaining) observed when solving GPQA questions. Comparison between a Llama model trained with curated pretraining data (orange) that amplified these behaviors in mathematical reasoning versus a behavior-minimized control model (blue).
RL training on Countdown transfers to other domains.
Next, we tested, a Qwen model trained with RL on Countdown, and the base Qwen model on 200 questions from GPQA and MATH each. The results (see Fig. 14) demonstrate a similar accentuation of learned behaviors, with the RL finetuned model demonstrating higher frequencies of all four cognitive behaviors. Surprisingly, we find a performance increase from 38% to 50% with our prompt format and 4-shot 44% to zero-shot 50% with Qwen’s 4-shot prompt format on the MATH dataset. Similar to the previous finding, we observe substantial increases in backtracking (0.15 to 0.59) and verification (0.02 to 0.29) in MATH and on GPQA, all four behaviors show meaningful increases: verification (0.008 to 0.03), backtracking (0.019 to 0.114), subgoal setting (0.478 to 0.616), and backward chaining (0.128 to 0.313). Notably, subgoal setting and backward chaining were not demonstrated by the Countdown RL finetuned Qwen model on Countdown (see Fig. 1 (Middle)) but despite this, these particular cognitive behaviors show accentuated activations post RL finetuning. These results taken together suggest that cognitive behaviors induced in the context of mathematical reasoning have a larger signature across other question-answering domains and beyond the training domain.
Figure 14: Transfer of Behaviors to Question Answering learned through RL. Average frequency of four cognitive behaviors (verifications, backtracking, subgoal setting, and backward chaining) observed when solving (a) GPQA and (b) MATH questions. Comparison between a Qwen model RL finetuned on Countdown (green) that amplified these behaviors in mathematical reasoning versus the base Qwen model (purple).
Appendix I Results for Instruction-tuned Models
We investigated how instruction tuning affects the emergence and amplification of cognitive behaviors with reinforcement learning (RL). Instruction tuning teaches models to reliably follow user instructions, but its impact on cognitive behaviors through RL is unclear. Toward this, we ran our behavioral probes on instruction-tuned variants of Llama-3.2-3B and Qwen-2.5-3B models on Countdown through RL training, following the same methodology applied to their base model counterparts. Our results (see Fig. 15) show that instruction-tuned models from both families nearly saturate task score and use more of their token budgets than the base models. Moreover, we observed the emergence and progressive strengthening of key cognitive behaviors—particularly verification and backtracking. These findings suggest that instruction-tuning datasets not only teach models to follow instructions but also contain examples of the cognitive behaviors that then surface during subsequent RL finetuning.
Figure 15: Analysis of the effect of instruction-tuning on the four cognitive behaviors. Plots compare (a) score and (b) response lengths between base models and their instruction tuned counterparts; and the four key cognitive behaviors through the course of RL on Countdown with (c) Qwen2.5-3B-Instruct and (d) Llama3.2-3B-Instruct.